An annuity immediate pays$1000 at the end of the first year. The payment increases by 3% per year to compensate for inflation. What is the present value of this annuity on the basis of a rate of 7%, if it runs for 20 years?
1
Expert's answer
2020-10-05T18:21:00-0400

solution

"Payment, P=1000"

"Growth\ rate, g=0.03"

"Interest,r=0.07"

"Period,n=20"



"Present\ value,PV=\frac{P}{r-g}[1-(\frac{1+g}{1+r})^n]"

"=\frac{1000}{0.07-0.03}[1-(\frac{1.03}{1.07})^{20}]=13331.6634"

Answer: the present value is $13,331.66